### v1

# Constructing bounded orbits of special types on homogeneous spaces

Abstract.
Let $X=G/\Gamma$ be a quotient of a real Lie group by a non-uniform lattice. Consider a one-parameter subgroup $F$ of $G$ that is $\operatorname{Ad}$-diagonalizable over $\mathbb{C}$} and whose action on $(X,m_{X})$ is mixing. In this dynamical system we study the set of points $x\in X$ with a precompact orbit, written as $E(F,\infty)$, which is known to be a dense subset of $X$ of full Hausdorff dimension. We prove that $E(F,\infty)$ is indecomposable in the following sense: given any $y\in E(F,\infty)$, the set of $x\in E(F,\infty)$ for which $y\in\overline{F_{+}x}$, where $F_{+}$ denotes the positive ray in $F$, is uncountable and dense in $E(F,\infty)$. When the dimension of the neutral subgroup of $G$ with respect to $F$ is $1$ we demonstrate, for any $\varepsilon>0$, the existence of many points $x\in X$ whose orbit closure $\overline{F_{+}x}\subset X$ is compact and has Hausdorff dimension at least $\dim X-\varepsilon$.
Key words and phrases:
homogeneous flows, orbit closures, bounded orbits, mixing
1991 Mathematics Subject Classification:
37A17; 37A25, 37D40, 11J70
The second-named author was supported by NSF grant DMS-2155111.
1. Introduction
Let $X$ be a metric space, and let $F$ be either of the (semi) groups $\mathbb{R}$} or $\mathbb{Z}$} or ${\mathbb{R}}_{\geq 0}$ or ${\mathbb{Z}}_{\geq 0}$ represented as a family of self-maps of $X$. We will write $F=\{g_{t}\}$. For $g\in F$ we will denote its action on $X$ by $(g,x)\mapsto gx$. Now fix a subset $Y\subset X$ and define the set of points of $X$ approaching $Y$ by $F$ as
$A(F,Y):=\big\{x\in X:\exists\text{ an unbounded sequence }(g_{t_{k}})_{k\in{\mathbb{N}}}\subset F\text{ with }{\lim_{k\to\infty}g_{t_{k}}x\in Y}\big\}.$
Its complement, the set of points of $X$ escaping $Y$ by $F$ will be denoted by E(F,Y)E(F,{Y}), namely
$E(F,{Y}):=\bigcap_{y\in Y}\big\{x\in X:y\notin\overline{(F\smallsetminus F_{0})x}\text{ for some bounded }F_{0}\subset F\big\}.$
If $X$ is not bounded, we make the convention that $x\in A(F,\infty)$ when the orbit $Fx$ is unbounded.
Take $y\in Y$ and consider $A(F,y):=A(F,\{y\})$ and $E(F,y):=E(F,\{y\})$. A natural question one could ask is how large the sets $A(F,y)$ and $E(F,y)$ can be. If the action admits an ergodic invariant probability measure, or has some hyperbolic behavior, some instances of this question can be answered. Namely, under the assumption of existence of an ergodic measure $\mu$ of full support the set $A(F,y)$ has full measure for any $y\in X\cup\{\infty\}$. Additionally, for many hyperbolic systems one can prove that the $\mu$-null sets $E(F,y)$ are thick, that is, their intersection with any nonempty open set has full Hausdorff dimension. For results of this kind see e.g. [D2] for toral endomorphisms, [U, Do] for Anosov flows, and [D1, KMa, Kl1, AGK] for partially hyperbolic flows on homogeneous space. Furthermore, in many cases these sets can be proved to be winning in the sense of Schmidt games (see e.g. [AGK, BFK, D1, D2, Ts1, Ts2, Wu1, Wu2]). Consequently in these cases it follows that both $A(F,y)$ and $E(F,y)$ are thick. Moreover, both full-measure and winning conditions are stable with respect to countable intersections; thus for any choice of countably many semigroups $F_{i}$ and points $y_{i}\in X\cup\{\infty\}$ for which the above conclusions can be established, it holds that both $\bigcap_{i}A(F_{i},y_{i})$ and $\bigcap_{i}E(F_{i},y_{i})$ are thick.
This paper addresses the following question: what if one considers a mixed case, that is, investigate the intersection
$A(F,y)\cap E(F,z),$} (1.1)
where $y,z\in X\cup\{\infty\}$? Problems of this type are amenable neither to the full-measure argument, nor to the technique based on Schmidt games. Also it is clear that some conditions on $y,z$ should be imposed; for example if $y=z$ it is clear that the intersection (1.1) is empty.
In what follows we restrict our attention to $F\cong{\mathbb{R}}$, also denoting $F_{+}:=\{g_{t}:t\geq 0\}$. Assume that the $F$-action on $X$ is continuous. Then let us record a natural obstruction to the set
$A(F_{+},y)\cap E(F_{+},z)$} (1.2)
being non-empty:
Lemma 1.1.
If $y\in X$ and $z\in X\cup\{\infty\}$ are such that $y\in A(F,z)$ (that is, $z$ can be approached by the trajectory of $y$ either in the positive or in the negative time direction), then
$A(F_{+},y)\subset A(F_{+},z).$
Proof.
Let $x\in A(F_{+},y)$, and let $(s_{\ell})_{\ell\in{\mathbb{N}}}$ be an unbounded sequence in ${\mathbb{R}}_{\geq 0}$ with
$\lim_{\ell\to\infty}g_{s_{\ell}}x=y.$
Let $W\subset X$ be an open neighborhood of $z$. In the case, when $z=\infty$, we take $W$ to be an open set with bounded complement.
By the hypothesis, there exists a time $t\in{\mathbb{R}}$ for which $g_{t}y\in W$. By the continuity of the action we have
$\lim_{\ell\to\infty}g_{s_{\ell}}g_{t}x=g_{t}y\in W.$
Note that the sequence $(s_{\ell}+t)_{\ell\in{\mathbb{N}}}$ is unbounded and eventually positive. Since $W$ was arbitrary, we conclude that $x\in A(F_{+},z)$. ∎
In view of the above lemma it makes sense to look for conditions guaranteeing that the set (1.2) is reasonably large under the assumption that $y\in E(F,z)$. In the present paper we will do it in the following set-up: we let $X=G/\Gamma$, where $G$ is a real Lie group and $\Gamma\subset G$ is a lattice, and consider the $\mathbb{R}$}-action on $X$ via the left multiplication by a one-parameter subgroup of $G$:
$\big\{g_{t}:=\exp(ta_{0})\in G:t\in{\mathbb{R}}\big\},$ (1.3)
where $a_{0}$} is an element of ${\mathfrak{g}}:=\operatorname{Lie}(G)$, the Lie algebra of $G$. We shall fix once and for all an inner product on $\mathfrak{g}$} to induce a right-invariant Riemannian metric $\operatorname{dist}_{G}$ on $G$. Note that the invariance of the metric implies that it is geodesically complete, and so the Hopf–Rinow theorem implies that it is proper (closed and bounded sets are compact). Then one obtains a metric on $X=G/\Gamma$ by setting
$\operatorname{dist}_{X}(g_{1}\Gamma,g_{2}\Gamma):=\inf\big\{\operatorname{dist}_{G}(g_{1}\gamma,g_{2}):\gamma\in\Gamma\big\}.$
This is the set-up in which we will attempt to study sets of the form (1.2). We will denote by $m_{X}$ the $G$-invariant Haar probability measure on $X$, and by $m_{G}$ the Haar measure on $G$ scaled so that the measure of any fundamental domain for $\Gamma$ is equal to $1$. Further, in this paper we will assume that $X$ is noncompact and $z=\infty$; that is, we will seek to construct bounded $F^{+}$-trajectories with prescribed limit points.
Recently the second- and third-named authors proved the following result. Take $G=\operatorname{SL}_{2}({\mathbb{R}})$, $\Gamma=\operatorname{SL}_{2}({\mathbb{Z}})$ and $X=G/\Gamma$. For $t,\alpha\in{\mathbb{R}}$ consider
$g_{t}:=\left[{\begin{array}[]{cc}e^{t}&0\\ 0&e^{-t}\end{array}}\right]\in G,\ F=\{g_{t}:t\in{\mathbb{R}}\},$
and
$x_{\alpha}:=\left[{\begin{array}[]{cc}1&\alpha\\ 0&1\end{array}}\right]\Gamma\in X.$}{}
Theorem 1.2.
[KR2, Theorem 1.3] Let $X$ and $F$ be as above, and let $y\in E(F,\infty)$; then
$\dim\big(\{\alpha\in{\mathbb{R}}:x_{\alpha}\in A(F_{+},y)\cap E(F_{+},\infty)\}\big)=1.$ (1.4)
Here and hereafter "dim" stands for Hausdorff dimension. In fact, even though it was not mentioned in [KR2], the proof actually demonstrates the thickness of the set in the left-hand side of (1.4). Also, by the standard "slicing" argument (see §2), it can be easily derived from Theorem 1.2 that for any $y\in E(F,\infty)$ the set $A(F_{+},y)\cap E(F_{+},\infty)$ is thick. The proof of the above theorem used the technique of continued fractions, and in fact the result can be interpreted within the framework of Diophantine approximation. Namely it constructs numbers that are badly approximable but not $\nu$-Dirichlet improvable, where $\nu$ is an irreducible norm on ${\mathbb{R}}^{2}$. See [KR1, KR2, KRS, AD] for more detail, and [MS] for some new results in that direction.
One of the main results of this paper is a generalization of Theorem 1.2 to a wider class of flows on homogeneous spaces.
Theorem 1.3.
Let $X=G/\Gamma$ and $F=\{g_{t}:t\in{\mathbb{R}}\}\subset G$, where
  • $G$ is a Lie group with discrete center,
  • $\Gamma$ is a non-uniform lattice in $G$,
  • $g_{t}$ is as in (1.3), where $a_{0}\in{\mathfrak{g}}\smallsetminus\{0\}$ is $\operatorname{ad}$}-diagonalizable over $\mathbb{C}$};
  • the action of $F$ on $(X,m_{X})$ is mixing.
Let $y\in E(F,\infty)$. Then for any non-empty open $U\subset X$ we have
$\dim\big(U\cap A(F_{+},y)\cap E(F_{+},\infty)\big)\geq\dim X-\dim Z+1,$ (1.5)
where $Z\subset G$ is the neutral subgroup of $G$ with respect to $F$ .
See Definition 2.2 below for a precise definition of $Z$. In view of Moore’s Theorem [Mo], the mixing assumption is satisfied when $G$ is a connected semi-simple Lie group with finite center and with no compact factors, and $\Gamma\subset G$ is irreducible. However mixing is not restricted to the semi-simple case, see e.g. [BM, Kl2]. Note that when $G=\operatorname{SL}_{2}({\mathbb{R}})$ (that is, under the assumptions of Theorem 1.2) we know that $\dim Z=\dim F=1$, and thus the set $A(F_{+},y)\cap E(F_{+},\infty)$ is thick. Moreover, in the latter case one can generalize Theorem 1.2 as follows.
Theorem 1.4.
Let $G$, $\Gamma$ and $F$ be as in Theorem 1.3, and assume in addition that $\dim Z=1$. Then for any compact F-invariant subset $B$ of $E(F,\infty)$ there exists a thick set of $x\in E(F_{+},\infty)$ such that the closure of the (bounded) trajectory $F_{+}x$ contains $B$.
In addition to $G=\operatorname{SL}_{2}({\mathbb{R}})$, the assumptions of the above theorem hold when $G=\operatorname{SL}_{2}({\mathbb{R}})\ltimes{\mathbb{R}}^{2}$; the choice $\Gamma=\operatorname{SL}_{2}({\mathbb{Z}})\ltimes{\mathbb{Z}}^{2}$ yields the space of two-dimensional unimodular grids, cf. [Kl2].
Corollary 1.5.
Let $G$, $\Gamma$, $X$ and $F$ be as in Theorem 1.4; then for any $\varepsilon>0$ there exists a thick set of $x\in X$ such that the orbit closure $\overline{F^{+}x}$ is compact and has Hausdorff dimension greater than $\dim X-\varepsilon$.
Proof.
When $X$ and $F$ are as in Theorem 1.3, it follows from [KMa, Theorem 1.1] that, given $\varepsilon>0$, one can find a large enough compact set $K\subset X$ such that the compact $F$-invariant set
$B=\{x\in X:Fx\subset K\}$
has Hausdorff dimension greater than $\dim X-\varepsilon$. Under the additional hypothesis of $\dim Z=1$, an application of Theorem 1.4 to $B$ then gives the result. ∎
Acknowledgements. The second-named author is grateful to ETH (Zürich) for its hospitality during a sabbatical stay in 2022, and to Barak Weiss for helpful discussions. The third-named author thanks N. Chandgotia, T. Mesikepp and Y. Peres for holding public-access office hours at TIFR-CAM, Peking University and Tsinghua University respectively.
2. Reduction to the expanding subgroup
In this and the remaining part of the paper we let $G$, $\Gamma$ and $F$ be as in Theorem 1.3. Recall that $F=\{g_{t}=\exp(ta_{0}):t\in{\mathbb{R}}\}$, where ${a_{0}}\in{\mathfrak{g}}\smallsetminus\{0\}$ is $\operatorname{ad}$}-diagonalizable over $\mathbb{C}$. Define
$\begin{split}{\mathfrak{h}}&:=\left\{b\in{\mathfrak{g}}:\lim_{t\to-\infty}\operatorname{Ad}(g_{t})b=0\right\}\\ {\mathfrak{n}}&:=\left\{b\in{\mathfrak{g}}:\sup_{t\in{\mathbb{R}}}\|\operatorname{Ad}(g_{t})b\|<\infty\right\}\\ {\mathfrak{h}}^{-}&:=\left\{b\in{\mathfrak{g}}:\lim_{t\to\infty}\operatorname{Ad}(g_{t})b=0\right\}.\end{split}$
From the diagonalizability assumption, we have
${\mathfrak{g}}={\mathfrak{h}}\oplus{\mathfrak{n}}\oplus{\mathfrak{h}}^{-}.$ (2.1)
In fact, if one considers the subalgebras of $\mathfrak{g}$$\otimes_{\mathbb{R}}{\mathbb{C}}$ which are direct sums of eigenspaces for ${\operatorname{ad}}_{a_{0}}\otimes 1$ corresponding to eigenvalues with positive, zero and negative real part respectively, then one obtains the above algebra decomposition (2.1) simply by intersecting with $\mathfrak{g}$}.
Remark 2.1.
Note that $\mathfrak{n}$} is a proper subalgebra of $\mathfrak{g}$}, otherwise $\operatorname{Ad}(g_{t})$ would be a compact group and the mixing hypothesis would fail. Moreover, since $G$ admits a lattice, the modular function of $G$ is constant. Thus $\operatorname{{\operatorname{ad}}_{a_{0}}}$ has trace 0, and, therefore, both $\mathfrak{h}$} and $\mathfrak{h}$$^{-}$ are nonzero.
Definition 2.2.
[Expanding, contracting and neutral subgroups] We let $H$, $H^{-}$ and $Z$ be the connected Lie subgroups of $G$ associated to the algebras $\mathfrak{h}$}, $\mathfrak{h}$$^{-}$ and $\mathfrak{n}$} respectively. We fix $m_{H}$, a Haar measure on $H$. We further consider the connected subgroup $\tilde{H}$ associated to the Lie algebra ${\mathfrak{n}}\oplus{\mathfrak{h}}^{-}$ and a left Haar measure $m_{\tilde{H}}$ on $\tilde{H}$, scaled so that $m_{G}$ is a direct product of $m_{H}$ and $m_{\tilde{H}}$, see [B, Ch. VII, §9, Proposition 13] or [Kn, Theorem 8.32] for justification.
By ordering the eigenspaces of ${\operatorname{ad}}_{a_{0}}\otimes 1$ with respect to their real parts, we see that $H$ is nilpotent and hence unimodular. Thus, $m_{H}$ is a bi-invariant Haar measure on $H$. For future reference, we also note that $H$ is simply connected. Indeed, note that $\operatorname{Ad}|_{H}:H\to\operatorname{Ad}(H)$ is a covering map. By choosing a basis compatible with the decomposition in (2.1), $\operatorname{Ad}(H)$ can be realized as a subgroup of the group of upper-triangular matrices. Thus $H$ covers a simply-connected space (see [Kn, Corollary 1.134]) and must itself be simply connected.
The inner product on $\mathfrak{g}$} induces a norm which we denote by $\|\cdot\|$. Further, using the induced inner product on $\mathfrak{h}$}, $\mathfrak{h}$$^{-}$ and $\mathfrak{n}$} we obtain right-invariant Riemannian distance functions on $H$, $H^{-}$ and $Z$ denoted by $\operatorname{dist}_{H}$, $\operatorname{dist}_{H^{-}}$ and $\operatorname{dist}_{Z}$. In view of the commutation relation $\operatorname{Ad}\big(\exp(b)\big)=e^{{\operatorname{ad}}_{b}}$ for all $b\in{\mathfrak{g}}$, we have
$g_{t}\exp(b)g_{-t}=\exp(e^{t{\operatorname{ad}}_{a_{0}}}b)\ \ \text{for all $b\in{\mathfrak{g}}$ and }t\in{\mathbb{R}}.$ (2.2)
In particular, $H,H^{-}$ and $Z$ are invariant under conjugation by $g_{t}$.
For $g\in G$, $x\in X$, $a\in{\mathfrak{g}}$ and $r>0$, we use the notation
$B_{G}(g,r),\ B_{X}(x,r)\text{ and }B_{\mathfrak{g}}(a,r)$
to denote balls in $G$ with respect to $\operatorname{dist}_{G}$, balls in $X$ with respect to $\operatorname{dist}_{X}$ and balls in $\mathfrak{g}$} with respect to the norm induced by the inner product. When the base-point is the identity in $G$ or the zero element in $\mathfrak{g}$}, we use the notation $B_{G}(r)$ or $B_{\mathfrak{g}}(r)$.
Definition 2.3 (Conjugation by $g_{t}$).
We use the notation, for each $t\in{\mathbb{R}}$,
$\Phi_{t}:G\to G;\ \Phi_{t}(g):=g_{t}gg_{-t}.$
In view of (2.2), $\Phi_{t}$ restricts to a smooth group isomorphism of the groups $H$, $H^{-}$ and $Z$. We can thus define the Radon–Nikodym derivative
$\frac{dm_{H}}{d(\Phi_{t})_{*}m_{H}}=e^{\chi t},\text{ where }\chi:={\text{tr}}\,{\operatorname{ad}}_{a_{0}}|_{{\mathfrak{h}}}.$ (2.3)
The equality above follows since $\exp:{\mathfrak{h}}\to H$ is a diffeomorphism and $m_{H}$ is the pushforward of a translation invariant measure on $\mathfrak{h}$} (see the proof of [Kn, Theorem 1.127]). Note that $\chi>0$.
The following lemma summarizes the expanding, contracting and neutral properties of $\Phi_{t}$ restricted to the subgroups $H$, $H^{-}$ and $Z$. It easily follows from Definition 2.2 and the locally bi-Lipschitz property of the exponential map ${\mathfrak{g}}\to G$.
Lemma 2.4.
If $h\in H$ and $h^{-}\in H^{-}$, we have
[…]
Also there exist constants $C,{r_{0}}>0$ such that
$\operatorname{dist}_{Z}\big({\Phi_{t}(z)},e_{G}\big)\leq C\operatorname{dist}_{Z}(z,e_{G})$}.
for all $z\in B_{Z}({r_{0}})$ and $t\in{\mathbb{R}}$.
The theorems in the previous section will be reduced to the following.
Theorem 2.5.
Let $X=G/\Gamma$ and $F=\{g_{t}\in G:t\in{\mathbb{R}}\}$ be as in Theorem 1.3. Let $\sigma_{0}>0$, $x_{0}\in X$, and let $B$ be a countable set with ${FB}\subset X$ bounded; in particular, $B\subset E(F,\infty)$. Then the set of $h\in H$ with the properties
  • $hx_{0}\in E(F_{+},\infty)$
  • for every $y\in B$, $hx_{0}\in A\big(F_{+},B_{Z}(\sigma_{0})y\big)$
is thick in $H$.
We now give the
Proof of Theorem 1.3 from Theorem 2.5.
Let $U\subset X$ be open, and let $y\in E(F,\infty)$. Fix $x_{0}\in U$, and let $r>0$ be such that the map
$\varphi:B_{Z}(r)\times B_{H^{-}}(r)\times B_{H}(r)\to U;\ \varphi(z,h^{-},h)=zh^{-}hx_{0}$
is a diffeomorphism onto an open subset of $U$. It follows from the definition of $\mathfrak{n}$} that
[…]
Take $r_{0},C$ as in Lemma 2.4. By shrinking $r$ if needed we can assume that $r<r_{0}$. Apply Theorem 2.5 with $\sigma_{0}=r/C$ and $B=\{y\}$. This gives us a set
$S:=\left\{h\in B_{H}(r):hx_{0}\in E(F_{+},\infty)\cap A\big(F_{+},B_{Z}(\sigma_{0})y\big)\right\}$
with $\dim S=\dim H$. Then for any $s\in{\mathbb{R}}$, $z\in Z$ and $h^{-}\in H^{-}$ compute
$\begin{split}g_{t}g_{s}zh^{-}hx_{0}&=g_{s}{\Phi_{t}(z)}\Phi_{t}(h^{-})g_{t}hx_{0}.\end{split}$ (2.5)
Lemma 2.4 thus shows that $g_{s}zh^{-}Sx_{0}\subset E(F_{+},\infty)$. Further, let $h\in S$, and suppose that $hx_{0}\in A(F_{+},zy)$ with $z\in B_{Z}({\sigma_{0}})$. This gives us an unbounded sequence $(t_{n})_{n\in{\mathbb{N}}}\subset{\mathbb{R}}_{\geq 0}$ with
$\lim_{n\to\infty}g_{t_{n}}hx_{0}=zy.$
Choose a subsequence $(t_{n_{k}})$ and $t_{0}$ as in (2.4), and let $w=\Phi_{t_{0}}^{-1}(z^{-1})$. We then see that for any $h^{-}\in H^{-}$ and $s\in{\mathbb{R}}$
$\lim_{k\to\infty}g_{t_{n_{k}}-s}\big(g_{s}wh^{-}h\big)x_{0}=\lim_{k\to\infty}\Phi_{t_{n_{k}}}(w)\Phi_{t_{k}}(h^{-})g_{t_{n_{k}}}hx_{0}=z^{-1}zy=y.$ (2.6)
The computations in (2.5) and (2.6) show the following. If $h\in S$, then there exists some $w\in B_{Z}(r)$ such that for all $s\in{\mathbb{R}}$ with $g_{s}w\in B_{Z}(r)$ and all $h^{-}\in B_{H^{-}}(r)$ we have
$g_{s}wh^{-}hx_{0}\in E(F_{+},\infty)\cap A(F_{+},y).$
Thus, if we define
$T:=\left\{(z,h^{-},h)\in B_{Z}(r)\times B_{H^{-}}(r)\times B_{H}(r):zh^{-}hx_{0}\in E(F_{+},\infty)\cap A(F_{+},y)\right\}$
and let $\pi:B_{Z}(r)\times B_{H^{-}}(r)\times B_{H}(r)\to B_{H}(r)$ be the projection onto the third factor, then in the commutative diagram
[…]
we have
$\dim\big(\pi^{-1}(h)\cap T\big)\geq 1+\dim H^{-}\text{ for all }h\in S.$
We now record the following dimension estimate.
Lemma 2.6 (Lemma 1.2.1 in [Kl1]).
Let $M_{1},M_{2}$ be Riemannian manifolds, and let $\pi:M_{1}\times M_{2}\to M_{2}$ be the natural projection. Let $E\subset M_{1}\times M_{2}$. Then
$\dim E\geq\dim\pi(E)+\inf_{p\in\pi(E)}\dim\left(\pi^{-1}(p)\cap E\right).$
Proof.
The use of a covering and a bi-Lipschitz equivalence reduces the statement to $M_{1}$ and $M_{2}$ being open subsets in Euclidean spaces, in which case the estimate is given by a theorem of Marstrand, see [Ma, Theorem 1] or [F, Theorem 5.8] [^1] In the aforementioned references the theorems is stated and proved for the case $\dim M_{1}=\dim M_{2}=1$, but the same proof works for arbitrary dimensions.. See also [EH, Theorem 1.1] for a modern treatment. ∎
Now let us apply the above lemma with $M_{2}=B_{H}(r)$, $M_{1}=B_{Z}(r)\times B_{H^{-}}(r)$ and $E=\pi^{-1}(S)\cap T$. Note that $\pi(E)=S$. We thus obtain
$\begin{split}\dim T&\geq\dim S+1+\dim H^{-}\\ &=\dim H+1+\dim H^{-}=\dim X-\dim Z+1.\end{split}$
Since $T$ is mapped to $U$ by $\varphi$, we have
$\dim\big(U\cap A(F_{+},y)\cap E(F_{+},\infty)\big)\geq\dim X-\dim Z+1,$
which completes the proof. ∎
We also have
Proof of Theorem 1.4 from Theorem 2.5.
By the hypothesis we have a compact, $F$-invariant subset $B\subset E(F,\infty)$. We are to construct points $x$ with $B\subset\overline{F_{+}x}$. If $B^{\prime}\subset B$ is countable with $\overline{B^{\prime}}=B$, then clearly it suffices to construct points $x$ with $B^{\prime}\subset\overline{F_{+}x}$. Thus, from the outset we assume that $B\subset E(F,\infty)$ is countable with $FB\subset X$ bounded.
Recall also that in this case we have ${\mathfrak{n}}={\mathbb{R}}a_{0}$ and $Z=F$. For any non-empty open $U\subset X$ we can choose a base point $x_{0}\in U$ and apply Theorem 2.5 with an arbitrary $\sigma_{0}>0$ and with $B\subset E(F,\infty)$ countable as above. This gives us a thick set of $h\in H$ such that, for every $y\in B$,
$hx_{0}\in E(F_{+},\infty)\cap A\big(F_{+},B_{Z}(\sigma_{0})y\big)=E(F_{+},\infty)\cap A(F_{+},y).$
The same computations as in the previous proof show that $E(F_{+},\infty)\cap A(F_{+},y)$ is stable under multiplication by $zh^{-}$ for arbitrary $z\in Z$ and $h^{-}\in H^{-}$. Thus Lemma 2.6 allows us to conclude that
$\dim\big(U\cap E(F_{+},\infty)\cap A(F_{+},y)\big)\geq\dim H+1+\dim H^{-}=\dim X,$
which finishes the proof. ∎
We can now focus on proving Theorem 2.5.
3. Equidistribution and tree-like sets on the expanding group
In what follows we fix the decomposition (2.1) of the Lie algebra of $G$ and will work with the subgroups $H,H^{-},Z$ as in Definition 2.2. A standard consequence of mixing of the $F$-action on $X$, dating back to the Ph.D. Thesis of Margulis [M], is the equidistribution of $g_{t}$-translates of $H$-orbits as $t\to\infty$. Let us quote a special case (one-sided inequality) stated in [Kl2].
Proposition 3.1 (Proposition 2.4 in [Kl2]).
Let $V\subset H$ be a bounded measurable set. Let $Q\subset X$ be a compact set with $m_{X}(\partial Q)=0$, let $L\subset X$ be another compact set, and let $\eta>0$. Then there exists $t_{0}:=t_{0}(V,Q,L,\eta)>0$ such that, for all $t>t_{0}$ and any base point $x\in L$, we have
$m_{H}\left(\{h\in V:g_{t}hx\in Q\}\right)>m_{H}(V)m_{X}(Q)-\eta.$
The above equidistirbution theorem was used in [KMa] to prove the existence of bounded $F_{+}$-trajectories on $X=G/\Gamma$. We are going to use the same principle to construct trajectories visiting small neighborhoods of points of $X$. For $x\in X$ denote by $\pi_{x}$ the map $G\to X$ given by $\pi_{x}(g):=gx$, and by $\rho(x)$ the injectivity radius of $x$:
${\rho(x):=}\,\sup\{r>0:\pi_{x}\text{ is injective on }B_{G}(r)\}.$
If $K$ is a subset of $X$, let us denote by $\rho(K)$ the injectivity radius of $K$:
$\rho(K):=\inf_{x\in K}\rho(x)=\sup\{r>0:\pi_{x}\text{ is injective on }B_{G}(r)\ \ \forall\,x\in K\};$
it is known that $\rho(K)>0$ if and only if $K$ is bounded.
Let us restate the mixing assumption on the FF–action on $X$ as follows.
Proposition 3.2 (Theorem 2.1.2 in [KMa]).
The $F$-action on $X$ is mixing if and only if for any two compact families of functions $\Xi,\Psi\subset L^{2}(m_{X})$ and for any $\eta>0$ there exists $t_{0}>0$ such that, for all $t>t_{0}$, $\varphi\in\Xi$ and $\psi\in\Psi$,
$\left|\int\varphi(g_{t}x)\psi(x)\,dm_{X}(x)-\left(\int\varphi\,dm_{X}\right)\left(\int\psi\,dm_{X}\right)\right|<\eta.$
The next proposition is a uniform version of Proposition 3.1 with $Q$ replaced by a compact family of balls in $X$ (compact in the sense of $L^{2}(m_{X})$).
Proposition 3.3.
Let $V\subset H$ be bounded measurable with $m_{H}(\partial V)=0$. Let $K\subset X$ be compact and let $0<\sigma<\rho(K)$. Let $L\subset X$ be another compact set, and let $\eta>0$. Then there exists $t_{0}:=t_{0}(V,K,\sigma,L,\eta)>0$ such that, for all $t>t_{0}$, base point $x\in L$ and center $y\in K$,
$m_{H}\left\{h\in V:g_{t}hx\in B_{X}(y,\sigma)\right\}>m_{H}(V)m_{G}\big(B_{G}(\sigma)\big)-\eta.$
Proof.
The proof is a modification of that of [Kl1, Proposition 2.4]. By partitioning $V$ into finitely many smaller subsets with boundary of measure zero and noting the invariance of $m_{H}$, one can without loss of generality assume that $e_{H}\in V$, $m_{H}(V)<1$ and $\textup{diam}(V)<\rho(L)$. The latter condition implies that there is an open ball $U\subset G$ centered at $e_{G}$ and containing $V$ such that for all $x\in L$ the map $\pi_{x}$ is injective on $U$, and hence
$m_{X}(Ux)=m_{G}(U).$
Similarly we have, for any $0<\alpha<1$ and $y\in K$,
$m_{X}\big(B_{X}(y,\alpha\sigma)\big)=m_{G}\big(B_{G}(\alpha\sigma)\big).$
Fix $\alpha<1$ such that
$m_{G}\big(B_{G}(\alpha\sigma)\big)>m_{G}\big(B_{G}(\sigma)\big)-\frac{\eta}{2}.$ (3.1)
Now let $\tilde{H}$ and $m_{\tilde{H}}$ be as in Definition 2.2. Choose a nonempty open ball $\tilde{V}\subset\tilde{H}$ such that $\tilde{V}V\subset U$ and such that for all $\tilde{v}\in\tilde{V}$ and $t>0$ we have
$\operatorname{dist}_{G}\big(e_{G},\Phi_{t}(\tilde{v})\big)<(1-\alpha)\sigma.$
Now for $y\in K$, $x\in L$ and $t>0$ denote
$W:=\left\{h\in V:g_{t}hx\in B_{X}(y,\sigma)\right\}\ \text{ and }\ U^{\prime}:=\tilde{V}V.$
Claim 3.3.1.
The set $U^{\prime}x\cap g_{t}^{-1}B_{X}(y,\alpha\sigma)$ is contained in $\tilde{V}Wx$.
Proof.
Let $\tilde{v}\in\tilde{V}$, $h\in V$ be such that $g_{t}\tilde{v}hx\in B_{X}(y,\alpha\sigma)$. We have
$\begin{split}\operatorname{dist}_{X}(g_{t}hx,y)&\leq\operatorname{dist}_{X}(g_{t}hx,\Phi_{t}(\tilde{v})g_{t}hx)+\operatorname{dist}_{X}(g_{t}\tilde{v}hx,y)\\ &<(1-\alpha)\sigma+\alpha\sigma=\sigma.\end{split}$
Thus $g_{t}hx\in B_{X}(y,\sigma)$. ∎
We now use Proposition 3.2 with
$\Xi:=\left\{\mathds{1}_{U^{\prime}x}:x\in L\right\},\ \Psi:=\{\mathds{1}_{B_{X}(y,\alpha\sigma}):y\in K\},$
and with $\eta$ replaced by $\eta m_{\tilde{H}}(\tilde{V})/2$. This gives $t_{0}>0$ such that, for all $t>t_{0}$, $x\in L$ and $y\in K$, we have
$\left|m_{X}\big(U^{\prime}x\cap g_{t}^{-1}B_{X}(y,\alpha\sigma)\big)-m_{G}(U^{\prime})m_{G}\big(B_{G}(\alpha\sigma)\big)\right|<\eta m_{\tilde{H}}(\tilde{V})/2.$
In particular, using the choice of $m_{\tilde{H}}$ and Claim 3.3.1, we compute
$\begin{split}m_{\tilde{H}}(\tilde{V})m_{H}(W)&=m_{X}(\tilde{V}Wx)\\ &\geq m_{X}\big(U^{\prime}x\cap g_{t}^{-1}B_{X}(y,\alpha\sigma)\big)\\ &\geq m_{G}(U^{\prime})m_{G}\big(B_{G}(\alpha\sigma)\big)-\eta m_{\tilde{H}}(\tilde{V})/2\\ &=m_{\tilde{H}}(\tilde{V})m_{H}(V)m_{G}\big(B_{G}(\alpha\sigma)\big)-\eta m_{\tilde{H}}(\tilde{V})/2.\end{split}$
[…]
$\begin{split}m_{H}(W)&\geq m_{H}(V)m_{G}\big(B_{G}(\alpha\sigma)\big)-\eta/2\\ &\geq m_{H}(V)m_{G}\big(B_{G}(\sigma)\big)-m_{H}(V)\eta/2-\eta/2\\ &\geq m_{H}(V)m_{G}\big(B_{G}(\sigma)\big)-\eta\end{split}$
which gives the desired result. ∎
Definition 3.4 (Tessellation domains in $H$).
Following [KMa], we say that a subset $V$ of a Lie group $H$ is a tessellation domain relative to a countable subset $\Lambda$ of $H$ if
  • (a)
    $m_{H}(\partial V)=0$;
  • (b)
    $V\gamma_{1}\cap V\gamma_{2}=\varnothing$ for different $\gamma_{1},\gamma_{2}\in\Lambda$;
  • (c)
    $H=\bigcup\limits_{\gamma\in\Lambda}{\overline{V}\gamma}.$
For example, the set $\{(x_{1},\dots,x_{k}):|x_{i}|<1/2\}\subset{\mathbb{R}}^{k}$ is a tessellation domain relative to ${\mathbb{Z}}^{k}\subset{\mathbb{R}}^{k}$. This can be generalized to tessellations of connected simply connected nilpotent Lie groups as follows.
Lemma 3.5 (Proposition 3.3 in [KMa]).
Any connected simply connected nilpotent Lie group $H$ admits a bounded tessellation domain. More precisely, there exists a basis $\{b_{i}\}$ for ${\mathfrak{h}}=\operatorname{Lie}(H)$ such that for every $r>0$, the set
$V_{r}:=\exp\Big(\Big\{\sum_{i}c_{i}b_{i}\in{\mathfrak{h}}:|c_{i}|<r\Big\}\Big)$ (3.2)
is a tessellation domain in $H$.
Another result from [KMa] estimates the boundary effects of partitioning the $\Phi_{t}$-image of $V$ into translates $V\gamma$:
Proposition 3.6 (Corollary 3.5 in [KMa]).
For any tessellation $(V,\Lambda)$ of $H$ and any $\eta>0$ there exists $T>0$ such that
$t\geq T\quad\Longrightarrow\quad\#\{\gamma\in\Lambda:V\gamma\subset\Phi_{t}(V)\}\geq e^{\chi t}\cdot(1-\eta)\,.$ (3.3)
Definition 3.7 (Strongly tree-like collections and Cantor sets).
We say a collection of subsets $\mathcal{E}\subset 2^{H}$ is strongly tree-like if:
  • Each $E\in\mathcal{E}$ is a compact set with nonempty interior.
  • We have a partition
    $\mathcal{E}=\bigcup_{k\in{\mathbb{Z}}_{\geq 0}}\mathcal{E}_{k}$
    with each $\mathcal{E}_{k}$ being finite and $\mathcal{E}_{0}$ being a singleton.
  • If $E^{\prime}\in\mathcal{E}_{k}$ with $k\in{\mathbb{N}}$, there is a unique $E\in\mathcal{E}_{k-1}$ with $E^{\prime}\subset E$.
  • For each $E\in\mathcal{E}_{k}$ with $k\in{\mathbb{Z}}_{\geq 0}$, there exists $E^{\prime}\in\mathcal{E}_{k+1}$ with $E^{\prime}\subset E$.
  • If $E_{1},E_{2}\in\mathcal{E}_{k}$ are distinct, then $m_{H}(E_{1}\cap E_{2})=0$.
  • If we define
    $d_{k}:=\sup\{\operatorname{diam}_{H}(E):E\in\mathcal{E}_{k}\},\text{ then }\lim_{k\to\infty}d_{k}=0.$ (3.4)
We write
$\cup\mathcal{E}_{k}:=\bigcup_{E\in\mathcal{E}_{k}}E$
and define the limit set of the collection to be $E_{\infty}:=\bigcap_{k\in{\mathbb{N}}}\cup\mathcal{E}_{k}$.
Associated to a tree-like collection $\mathcal{E}$ we also define, for $E\in\mathcal{E}_{k}$,
$\operatorname{density}(\mathcal{E}_{k+1},E):=\frac{m_{H}((\cup\mathcal{E}_{k+1})\cap E)}{m_{H}(E)}.$
Further, for each $k\in{\mathbb{Z}}_{\geq 0}$, we set
$\Delta_{k}:=\inf\left\{\operatorname{density}(\mathcal{E}_{{k+1}},E):E\in\mathcal{E}_{k}\right\}$ (3.5)
Theorem 3.8 (Lemma 2.1 in [U]).
For a tree-like collection $\mathcal{E}$ and the resulting limit set $E_{\infty}$, we have the dimension estimate
$\dim H-\dim E_{\infty}\leq\limsup_{k\to\infty}\frac{\sum_{j=0}^{k}\log\Delta_{j}}{\log d_{k}}.$
4. The generic and special steps of the induction
The desired orbits in Theorem 2.5 are constructed inductively. We start with a base point $x_{0}$ and a tessellation domain $V\subset H$; the latter will always be chosen of the form $V=V_{r}$ as in (3.2), where $r>0$ is small enough. Then, roughly speaking, we choose a collection of subsets $\{W\subset V\}$ such that $g_{t}$-images of $Wx_{0}$ have a certain behavior. The expansive properties of the flow ensure that each $g_{t}Wx_{0}=\Phi_{t}(W)g_{t}x_{0}$ looks precisely like $Vx_{1}$ for some $x_{1}\in X$, and from there we begin the process all over again. The two types of behavior to be considered are 1) being bounded within a certain compact set, and 2) approaching a prescribed point $y$. We now formalize these two steps as follows.
Proposition 4.1 (The generic step).
Let $K\subset X$ be a compact set with $m_{X}(\partial{K})=0$ and let $\eta>0$. Then one can choose $\delta>0$ with the following property: for any $r<\delta$ there exist a tessellation $(V_{r},\Lambda_{r})$ of $H$ and a time $t_{0}>0$ such that for all points $x\in K$ and $t\geq t_{0}$ we have
$\left\{\gamma\in\Lambda_{r}:V_{r}\gamma\subset\Phi_{t}(V_{r})\text{ and }V_{r}\gamma g_{t}x\subset K\right\}\geq e^{\chi t}\big(m_{X}(K)-\eta\big)$}, (4.1)
where $\chi$ is as in (2.3).
Proof.
This is a simplified version of [KMa, Proposition 3.6]. We repeat the argument here for the sake of clarity. The proposition is true if $m_{X}(K)=0$ by choosing $\delta$ and $t_{0}$ arbitrarily. Thus we can assume that the interior of $K$ has positive measure.
Given $\eta>0$, we choose a compact subset $K^{\prime}\subset K$ and $\delta>0$ so that:
  1. (a)
    $m_{X}(\partial K^{\prime})=0$ and $m_{X}(K^{\prime})>m_{X}(K)-\eta/3$;
  2. (b)
    $V_{\delta}V_{\delta}^{-1}K^{\prime}\subset K$ (hence $V_{r}V_{r}^{-1}K^{\prime}\subset K$ for any $0<r<\delta$).
For any $r$ as above let $V=V_{r}$ and choose $\Lambda=\Lambda_{r}$ such that $(V,\Lambda)$ is a tessellation of $H$. Then for all $x\in X$ and $t>0$ we have the following estimate:
[…]
(the last step follows from (2.3) and Definition 2.3). Now apply Proposition 3.1 with $Q$ replaced by $K^{\prime}$, $L$ replaced by $K$, and $\eta$ replaced by $\frac{\eta m_{H}(V)}{3}$. This gives us a time $t_{1}>0$ such that, for all $x\in K$ and $t\geq t_{1}$,
$m_{H}\big(\left\{h\in V:g_{t}hx\in{K^{\prime}}\right\}\big)>m_{H}(V)\big(m_{X}({K^{\prime}})-\eta/{3}\big)$}. (4.3)
On the other hand, Proposition 3.6 yields a time $t_{2}>0$ such that, for all $t\geq t_{2}$,
$\left\{\gamma\in\Lambda:V\gamma\subset\Phi_{t}(V)\right\}\geq e^{\chi t}\left(1-{\eta}/{3}\right).$ (4.4)
Letting $t_{0}:=\max\{t_{1},t_{2}\}$ and combining (4.2), (4.3) and (4.4), for any $t>t_{0}$ we obtain
$\begin{split}\#\left\{\gamma\in\Lambda:V\gamma\subset\Phi_{t}(V)\text{ and }V\gamma g_{t}x\subset K\right\}&\geq e^{\chi t}(1-\eta/3)-e^{\chi t}\big(1-m_{X}(K^{\prime})+\eta/3\big)\\ &=e^{\chi t}\big(m_{X}(K^{\prime})-2\eta/3\big)\geq e^{\chi t}\left(m_{X}(K)-\eta\right),\end{split}$
which is the desired conclusion. ∎
We now come to the special step of the procedure which ensures that our constructed orbits have prescribed limit points.
Proposition 4.2 (The approach step).
For any compact $K\subset X$ there exists $\delta>0$ such that for any $r,\sigma<\delta$ one has the tessellation $(V_{r},\Lambda_{r})$ and a time $t_{0}>0$ such that the following holds. For any $t\geq t_{0}$ and $\varepsilon>0$ one can find $s_{\varepsilon}>s^{\prime}_{\varepsilon}>t$ such that for any $x\in K$ and any $y\in K$ with
$Fy\subset K\text{ and }\operatorname{dist}_{X}(Fy,\partial K)>1$} (4.5)
there exists $\gamma=\gamma(x,y,\varepsilon)\in\Lambda_{r}$ satisfying
  • (i)
    The set $W:=\Phi_{-s_{\varepsilon}}(V_{r}\gamma)$ is a subset of $V_{r}$.
  • (ii)
    For all $s\in[t,s_{\varepsilon}]$ we have $g_{s}Wx\subset K$.
  • (iii)
    For all $h\in W$, $\operatorname{dist}_{X}\big(g_{s^{\prime}_{\varepsilon}}hx,B_{Z}(\sigma)y\big)<\varepsilon$.
Proof of Proposition 4.2.
Take $\delta_{1}>0$ such that the multiplication map from $H\times H^{-}\times Z$ to $G$ maps onto $B_{G}(\delta_{1})$ and has an inverse on $B_{G}(\delta_{1})$. That is, we have a smooth and bi-Lipschitz (with respect to the supremum metric of the product space) map
$B_{G}(\delta_{1})\to H\times H^{-}\times Z$
which composes with multiplication to give the identity map on $B_{G}(\delta_{1})$. In particular, there exists a constant $c\in(0,1)$ such that, for all $r\leq\delta_{1}$,
$g\in B_{G}(cr)\implies g=hh^{-}z\text{ for some }(h,h^{-},z)\in B_{H}(r)\times B_{H^{-}}(r)\times B_{Z}(r).$
We may shrink $\delta_{1}$ further to ensure that
  • $\delta_{1}<1/3$;
  • $\operatorname{diam}_{X}(V_{\delta_{1}})<1/3$;
  • $\Phi_{t}\big(B_{H-}(\delta_{1})\big)\subset B_{G}(1/3)\text{ for all }t\geq 0$;
  • $\Phi_{t}\big(V_{\delta_{1}}(\overline{V_{\delta_{1}}})^{-1}\big)\subset B_{G}(1/3)$ for all $t\leq 0$;
  • $\Phi_{t}\big(B_{Z}(\delta_{1})\big)\subset B_{G}(1/3)$ for all $t\in{\mathbb{R}}$.
Now fix $\delta<\min\{\rho(K),\delta_{1}\}$. Thus, for every point $x\in K$ and $r<\delta$ the map
$B_{H}(r)\times B_{H^{-}}(r)\times B_{Z}(r)\to X;\ (h,h^{-},z)\mapsto hh^{-}zx$
maps onto a neighborhood of $B_{X}(x,cr)$.
Fix $r,\sigma<\delta$ and consider the tessellation domain $V=V_{r}$ from Lemma 3.5 with a corresponding discrete set $\Lambda=\Lambda_{r}$. Apply Proposition 3.3 with parameters $V$, $K$, $\sigma$, $L$ and $\varepsilon$ there replaced by $V_{r/2}$, $K$, $c\sigma$, $K$ and
$\frac{m_{H}(V_{r/2})m_{G}(B(c\sigma))}{2}$
respectively. From this we obtain a $t_{0}>0$ such that, for all $t>t_{0}$, $x\in K$, $y\in K$,
$m_{H}\left(\left\{h\in V_{r/2}:g_{t}hx\in B_{X}(y,c\sigma)\right\}\right)>\frac{m_{H}(V_{r/2})m_{G}(B(c\sigma))}{2}.$ (4.6)
We further assume $t_{0}$ is large enough so that
$\Phi_{-t}\left((B_{H}(\sigma))^{-1}\right)V_{r/2}\subset V_{3r/4}\text{ for all }t\geq t_{0}.$ (4.7)
We now fix $t\geq t_{0}$ and $\varepsilon>0$, choose $s_{1}>0$ large enough so that
$\Phi_{s_{1}}\big(B_{H^{-}}(\sigma)\big)\subset B_{H^{-}}(\varepsilon/2),$ (4.8)
and then choose $s_{\varepsilon}>s^{\prime}_{\varepsilon}:=s_{1}+t$ so that
$\Phi_{-s_{\varepsilon}}\big(V(\overline{V})^{-1}\big)V_{3r/4}\subset V$ (4.9)
and
$\Phi_{s^{\prime}_{\varepsilon}-{s_{\varepsilon}}}\big(V(\overline{V})^{-1}\big)\subset B_{H}(\varepsilon/2).$ (4.10)
Now fix $x\in K$ and $y\in K$ such that (4.5) holds. Inequality (4.6) applied with $y$ replaced by $g_{-s_{1}}y$, which is also contained in $K$, gives us an $h_{1}\in V_{r/2}$ with
$g_{t}h_{1}x\in B_{X}(g_{-s_{1}}y,c\sigma).$
The assumption on $\delta$ allows us to write $g_{t}h_{1}x$ as a product.
$g_{t}h_{1}x=h_{2}h^{-}zg_{-s_{1}}y\text{ for }h_{2}\in B_{H}(\sigma),\ h^{-}\in B_{H^{-}}(\sigma),\ z\in B_{Z}(\sigma).$ (4.11)
In particular, we have
$g_{t}\Phi_{-t}(h_{2}^{-1})h_{1}x=h^{-}zg_{-s_{1}}y.$ (4.12)
The stipulation in (4.7) ensures that we can write
$\Phi_{-t}(h_{2}^{-1})h_{1}x=h_{3}x\text{ where }h_{3}\in V_{3r/4}.$ (4.13)
Combining (4.12) and (4.13), we have
$g_{t}h_{3}x=h^{-}zg_{-s_{1}}y.$ (4.14)
Choose and fix $\gamma\in\Lambda$ with $\Phi_{s_{\varepsilon}}(h_{3})\in\overline{V}\gamma$. We write
$h_{3}=\Phi_{-s_{\varepsilon}}(h_{4}\gamma)\text{ for some }h_{4}\in\overline{V},$ (4.15)
which gives
$W:=\Phi_{-s_{\varepsilon}}(V\gamma)=\Phi_{-s_{\varepsilon}}(Vh_{4}^{-1})h_{3}.$
Since $h_{3}\in V_{3/4}$ and since we have the containment (4.9), we see that $W\subset V$. This proves Assertion (i) of the Proposition. We now compute, for $t\leq s\leq s_{\varepsilon}$,
[…]
where we have used the stipulations on $\delta$ and (4.5) for the last two containments. This proves Assertion (ii).
Finally we compute, as in (4.16),
[…]
where (4.10) and (4.8) are used to justify the last containment. This proves Assertion (iii) and finishes the proof of the proposition. ∎
5. Proof of Theorem 2.5
In this section our goal is to prove Theorem 2.5, that is, for given $x_{0}\in X$, construct many points of the form $hx_{0}$ such that their $F_{+}$-orbits are bounded, and, in addition, a given countable subset of $X$ is in their orbit closure. Our proof is a modification of a well-known scheme of constructing points with bounded orbits due to the second-named author and Margulis [KMa], see also [Kl1, Kl2]. Before dealing with the general case it will be instructive to briefly describe this construction in a special case.
Proof of Theorem 2.5 in the special case $B=\varnothing$.
Take $x_{0}\in X$ and a nonempty open subset $U$ of $H$. We need to prove that the Hausdorff dimension of the set
$\{h\in U:F_{+}hx_{0}\text{ is bounded}\}$} (5.1)
is equal to $\dim H$. Replacing $x_{0}$ by $hx_{0}$ for some $h\in U$ we can assume that $U$ is a neighborhood of identity in $H$.
Pick a compact set $K\subset X$ with $m_{X}(\partial K)=0$ containing $x_{0}$, and let
$\eta:=m_{X}(K)-m_{X}(K)^{2}.$} (5.2)
Then take $\delta>0$ furnished by Proposition 4.1 and choose $0<r<\delta$ such that $V_{r}\subset U$. By shrinking $r$ further we can assume that
$\operatorname{diam}_{H}(V_{r})<1/2.$} (5.3)
As before we denote $V=V_{r}$ and let $\Lambda=\Lambda_{r}$ be a corresponding discrete subset of $H$. The proposition gives us the value $t_{0}>0$ such that (4.1) holds for all $x\in K$ and $t\geq t_{0}$. Fix $t>t_{0}$ and let $\Lambda(x)$ denote the set appearing in the left hand side of (4.1), i.e.
$\Lambda(x):=\left\{\gamma\in\Lambda:V\gamma\subset\Phi_{t}(V)\text{ and }V\gamma g_{t}x\subset K\right\}.$} (5.4)
To estimate the Hausdorff dimension of the set (5.1) from below, for each $x\in K$ we are going to construct a strongly tree-like collection $\mathcal{E}$ of subsets of $H$ such that its limit set
$\mathcal{E}_{\infty}:=\bigcap_{k\in{\mathbb{N}}}\cup\mathcal{E}_{k}$
is contained in the set (5.1). We first let $\mathcal{E}_{0}:=\{\overline{V}\}$. Then define
$\mathcal{E}_{1}:=\left\{\Phi_{-t}(\overline{V}\gamma):\gamma\in\Lambda(x_{0})\right\}.$
That is, elements of $\mathcal{E}_{1}$ are of the form $E=\Phi_{-t}(\overline{V})h$ for some $h\in H$, chosen so that $E\subset\overline{V}$ and $g_{t}Ex_{0}\subset K$. Similarly, elements of $\mathcal{E}_{k}$ will be of the form
$E=\Phi_{-kt}(\overline{V})h\subset\overline{V}\text{ for some }h\in H,$
and it will follow from our inductive construction that
$E\in\mathcal{E}_{k}\Longrightarrow g_{it}Ex_{0}\subset K\text{ for }i=1,\dots,k.$} (5.5)
Namely, if $\mathcal{E}_{k}$ is chosen and we take $E\in\mathcal{E}_{k}$ such that
$E=\Phi_{-kt}(\overline{V})h\subset\overline{V}\text{ and }g_{it}Ex_{0}\subset K\text{ for }i=1,\dots,k,$
define the subcollection
$\mathcal{E}_{k+1}(E):=\left\{\Phi_{-(k+1)t}(\overline{V}\gamma):\gamma\in\Lambda({g_{kt}hx_{0}})\right\},$} (5.6)
and let $\mathcal{E}_{k+1}:=\cup_{E\in\mathcal{E}_{k}}\mathcal{E}_{k+1}(E)$. Then $\mathcal{E}_{k+1}(E)$ consists precisely of elements of $\mathcal{E}_{k+1}$ contained in $E$ (i.e. "children" of $E$), and it follows from the definition of $\Lambda({g_{kt}hx_{0}})$ that $g_{(k+1)t}Ex_{0}\subset K$ for any $E\in\mathcal{E}_{k+1}$.
The first five parts of Definition 3.7 follow readily from the construction and from $(V,\Lambda)$ being a tessellation of $H$. Also, if we define
$\lambda_{\text{min}}:=\min\{\Re\lambda:\lambda\text{ is an eigenvalue for }{\operatorname{ad}}_{a_{0}}\otimes 1\text{ with }\Re\lambda>0\},$
then a computation using the exponential map shows that for all $t$ sufficiently large,
$\operatorname{dist}_{H}\big(\Phi_{-t}(h),e_{G}\big)\leq e^{-\frac{\lambda_{\text{min}}t}{2}}\operatorname{dist}_{H}(h,e_{G})\text{ for all }h\in V.$ (5.7)
Note that the factor $1/2$ in the exponent is used to get rid of the constants that occur when comparing distances on $H$ with norms in $\mathfrak{g}$}. By increasing $t$ we can assume that (5.7) is satisfied. From (5.6) and the right-invariance of the metric we can see that for any $x\in K$ the diameter of any $E\in\mathcal{E}_{k}$ is equal to $\operatorname{diam}_{H}\big(\Phi_{-kt}(V)\big)$. Thus it follows from (5.3) and (5.7) that
$d_{k}=\sup\{\operatorname{diam}_{H}(E):E\in\mathcal{E}_{k}\}\leq e^{-\frac{\lambda_{\operatorname{min}}kt}{2}}.$ (5.8)
Hence (3.4) holds, and $E$ is a strongly tree-like collection. Clearly (5.5) implies that $g_{kt}E_{\infty}x_{0}\subset K$ for all $k\in{\mathbb{N}}$, hence $F_{+}E_{\infty}x_{0}$ is contained in $g_{[-t,0]}K$.
It remains to estimate the Hausdorff dimension of $E_{\infty}$ from below. Using (5.6) one can easily see that for any $k=0,1,\dots$ and for $E\in\mathcal{E}_{k}$ of the form $E=\Phi_{-kt}(\overline{V})h$,
$\operatorname{density}(\mathcal{E}_{k+1},E)=\frac{m_{H}\big(\cup\mathcal{E}_{k+1}(E)\big)}{m_{H}(E)}\geq e^{-\chi t}\#\Lambda({g_{kt}hx_{0}}).$
Thus Proposition 4.1 and (5.2) imply that for any $k=0,1,\dots$ one has
$\Delta_{k}=\inf\left\{\operatorname{density}(\mathcal{E}_{{k+1}},E):E\in\mathcal{E}_{k}\right\}\geq m_{X}(K)-\eta=m_{X}(K)^{2},$ (5.9)
and an application of Theorem 3.8 yields
$\dim H-\dim{E}_{\infty}\leq$ $\limsup_{k\to\infty}\frac{\sum_{j=0}^{k}\log\Delta_{j}}{\log d_{k}}$
[…]
The right hand side of the above inequality can be made arbitrary small by increasing $t$, which implies that the Hausdorff dimension of the set (5.1) is equal to $\dim H$. ∎
Now let us show how the above scheme can be modified to produce a proof of the general case of Theorem 2.5. The idea is to perform lots of generic steps as before, but occasionally insert approach steps.
Proof of Theorem 2.5 when $B\neq\varnothing$.
Fix a surjection $\beta:{\mathbb{N}}\to B$ such that $\beta^{-1}(y)$ is infinite for every $y\in B$, and a sequence of positive real numbers $(\varepsilon_{n})_{n\in{\mathbb{N}}}$ converging to zero.
Take a base-point $x_{0}\in X$. As before, it suffices to choose a neighborhood $U$ of the identity in $H$ and show that the set of $h\in U$ such that $hx_{0}$ lies in
$E(F_{+},\infty)\cap\bigcap_{y\in B}A(F_{+},B_{Z}(\sigma_{0})y)$ (5.10)
has full dimension in $H$. Choose a compact set $K$ with
$x_{0}\in\text{int}(K),\ m_{X}(\partial K)=0,\ FB\subset K\text{ and }\operatorname{dist}_{X}(FB,\partial K)>1.$
Apply Proposition 4.1 with this $K$ and $\eta=m_{X}(K)-m_{X}(K)^{2}$. This gives us a $\delta_{1}>0$. Now we may further apply Proposition 4.2 to $K$ to obtain another $\delta_{2}>0$. We fix
[…]
and obtain a tessellation domain $(V_{r},\Lambda_{r})$, which we write as $(V,\Lambda)$, and $t_{0}>0$ such that the conclusions of Propositions 4.1 and 4.2 hold. Fix $t>t_{0}$, and, as in the proof of the special case, by increasing $t$ if needed assume that (5.7) is satisfied. This way:
  • for any $x\in K$ we can define $\Lambda(x)\subset\Lambda$ as in (5.4) with $\Lambda(x)\geq e^{\chi t}m_{X}(K)^{2}$;
  • for any $\varepsilon>0$ one can choose $s_{\varepsilon}>s^{\prime}_{\varepsilon}>t$ such that any $x\in K$ and $y\in B$ come with $\gamma=\gamma(x,y,\varepsilon)\in\Lambda$ satisfying conditions (i)–(iii) of Propositions 4.2.
We now formalize the process of alternating generic and approach steps. Namely fix an injection $f:{\mathbb{N}}\to{\mathbb{N}}$ to be determined later; our strategy will consist of performing approach steps whenever $k$ is of the form $f(n)$, and generic steps otherwise. Specifically, for each $x\in K$ and $k\in{\mathbb{N}}$ we define a subset $\Lambda(x,k)$ of $\Lambda$ by
[…]
We also define a sequence of times $(t_{k})_{k\in{\mathbb{N}}}$ by
$t_{k}:=\begin{cases}t&\text{if }k\notin f({\mathbb{N}}),\\ s_{\varepsilon_{n}}&\text{if }k=f(n),\end{cases}$
and set $T_{0}=0$ and $T_{k}:=\sum_{j=1}^{k}t_{j}\text{ for each }k\in{\mathbb{N}}$, or, equivalently,
$T_{k}=\big(k-\#f^{-1}(\{1,\dots,k\})\big)t+\sum_{f(n)\leq k}s_{\varepsilon_{n}}=kt+\sum_{f(n)\leq k}(s_{\varepsilon_{n}}-t).$ (5.12)
We now come to the inductive construction. Define
$\mathcal{E}_{0}:=\{\overline{V}\},\ p_{0}:\mathcal{E}_{0}\to H;\ \overline{V}\mapsto e_{H}.$
Assume we have defined $\mathcal{E}_{j}$ and $p_{j}:\mathcal{E}_{j}\to H$ for $j=0,1,\dots,k$ with the properties that
  1. (a)
    Each $E\in\mathcal{E}_{j}$ is of the form
    $\Phi_{-T_{j}}(\overline{V})p_{j}(E)\ \text{ with }\ g_{T_{j}}p_{j}(E)x_{0}\in K.$
  2. (b)
    If $E\in\mathcal{E}_{j}$, then
    $g_{[0,T_{j}]}Ex_{0}\subset g_{[-t,t]}K.$
  3. (c)
    For any $j\in\{1,\dots,k\}$ with $j=f(n)$ for some $n\in{\mathbb{N}}$ there exists $T\in[T_{j-1},T_{j}]$ such that, for all $E\in\mathcal{E}_{j}$,
    $g_{T}Ex_{0}\subset B_{Z}(\sigma)B_{X}\big(\beta(n),\varepsilon_{n}\big).$
We then proceed to define, for each $E\in\mathcal{E}_{k}$, the collection of children $\mathcal{E}_{k+1}(E)$ as
$\mathcal{E}_{k+1}(E)$= $\text{ the set of all }E^{\prime}=\Phi_{-T_{k+1}}(\overline{V}\gamma)p_{k}(E),$ (5.13)
$\text{ where }\gamma\in\Lambda(g_{T_{k}}p_{k}(E)x_{0},k+1).$
We further define, for each $E^{\prime}$ as above,
$p_{k+1}(E^{\prime})=\Phi_{-T_{k+1}}(\gamma)p_{k}(E).$
Setting $\mathcal{E}_{k+1}:=\bigcup_{E\in\mathcal{E}_{k}}\mathcal{E}_{k}(E)$, we now proceed to verify the conditions (a)–(c) of the induction. For (a), note that each $E^{\prime}\in\mathcal{E}_{k+1}(E)$ is of the form $\Phi_{-T_{k+1}}(\overline{V})p_{k+1}(E^{\prime})$, and we must show that $g_{T_{k+1}}p_{k+1}(E^{\prime})x_{0}\in K$. We compute
$g_{T_{k+1}}p_{k+1}(E^{\prime})x_{0}=g_{T_{k+1}}\Phi_{-T_{k+1}}(\gamma)p_{k}(E)x_{0}=\left(\gamma g_{t_{k+1}}\right)g_{T_{k}}p_{k}(E)x_{0},$
which is contained in $K$ either from the definition of the set in (4.1), or from Proposition 4.2(ii). (Note that $V$ contains the identity.)
We come to (b). First note that each $E^{\prime}\in\mathcal{E}_{k+1}(E)$ as in (5.13) is a subset of $E$:
$E^{\prime}=\Phi_{-T_{k+1}}(\overline{V}\gamma)p_{k}(E)=\Phi_{-T_{k}}\circ\Phi_{-t_{k+1}}(\overline{V}\gamma)p_{k}(E)\subset\Phi_{-T_{k}}(V)p_{k}(E)=E$
where the containment holds from the either Proposition 4.1 or Proposition 4.2(i). Thus, it suffices to show $g_{T_{k}+s}E^{\prime}x_{0}\subset g_{[-t,t]}K$ for all $0\leq s\leq t_{k+1}$. For such an $s$, we compute
$g_{T_{k}+s}E^{\prime}x_{0}=g_{T_{k}+s}\Phi_{-T_{k+1}}(\overline{V}\gamma)p_{k}(E)x_{0}=g_{s}\Phi_{-t_{k+1}}(\overline{V}\gamma)g_{T_{k}}p_{k}(E)x_{0}.$
Now if $k+1\notin f({\mathbb{N}})$, we know that $t_{k+1}=t$, and the conclusion follows from Proposition 4.1. Otherwise, in view of (5.11), (5.13) and Proposition 4.2, for some $n\in{\mathbb{N}}$ we have
[…]
Hence the conclusion follows from Proposition 4.2(ii). Finally, to demonstrate (c) it suffices to only consider the approach step case. Then one can use (5.14) and Proposition 4.2(iii) to conclude that
$g_{T_{k}+s^{\prime}_{\varepsilon_{n}}}{E^{\prime}}x_{0}=g_{s^{\prime}_{\varepsilon_{n}}}\overline{W}g_{T_{k}}p_{k}(E)x_{0}$
lies in the $\varepsilon_{n}$-neighborhood of $B_{Z}(\sigma)\beta(n)$.
This completes the inductive construction. As before, it is easy to show that the family of sets $\mathcal{E}_{k}$ satisfies the requirements of Definition 3.7; in particular, in view of (5.3) and (5.7),
$d_{k}=\sup\{\operatorname{diam}_{H}(E):E\in\mathcal{E}_{k}\}\leq e^{\frac{-\lambda_{\operatorname{min}}T_{k}}{2}}.$ (5.15)
In light of (b), (c) and the infinitude of the fibers of $\beta$, the set in (5.10) contains $E_{\infty}x_{0}$.
It thus remains to estimate the Hausdorff dimension of $E_{\infty}$. Clearly
$\Delta_{k}=\inf\{\operatorname{density}(\mathcal{E}_{k+1},E):E\in\mathcal{E}_{k}\}$
depends on whether $k+1$ corresponds to a generic or a special step. In the former case estimate (5.9) still holds. Otherwise $k+1=f(n)$ for some $n\in{\mathbb{N}}$, $\mathcal{E}_{k+1}(E)$ is a singleton for every $E\in\mathcal{E}_{k}$, and $\Delta_{k}=e^{-\chi s_{\varepsilon_{n}}}$. Hence, analogously to (5.12), we can write
$\sum_{j=0}^{k}\log\Delta_{j}$ $\geq\big(k+1-\#f^{-1}(\{1,\dots,k+1\})\big)\log\big(m_{X}(K)^{2}\big)-\chi\sum_{f(n)\leq k+1}s_{\varepsilon_{n}}$ (5.16)
$=2(k+1)\log m_{X}(K)-\sum_{f(n)\leq k+1}\big(\chi s_{\varepsilon_{n}}+2\log m_{X}(K)\big).$
Now, using (5.12), (5.15) and (5.16) we can apply Theorem 3.8 to obtain
$\dim H-\dim{E}_{\infty}\leq$ $\limsup_{k\to\infty}\frac{\sum_{j=0}^{k}\log\Delta_{j}}{\log d_{k}}$
$\leq$ $\limsup_{k\to\infty}\frac{2(k+1)\log m_{X}(K)-\sum_{f(n)\leq k+1}\big(\chi s_{\varepsilon_{n}}+2\log m_{X}(K)\big)}{-\frac{1}{2}{\lambda_{\operatorname{min}}\big(kt+\sum_{f(n)\leq k}(s_{\varepsilon_{n}}-t)\big)}}$
$=$ $\limsup_{k\to\infty}\frac{2(k+1)\log\frac{1}{m_{X}(K)}+\sum_{f(n)\leq k+1}\big(\chi s_{\varepsilon_{n}}-2\log\frac{1}{m_{X}(K)}\big)}{\frac{1}{2}{\lambda_{\operatorname{min}}\big(kt+\sum_{f(n)\leq k}(s_{\varepsilon_{n}}-t)\big)}}$
$\leq$ $\frac{4\log\frac{1}{m_{X}(K)}}{\lambda_{\operatorname{min}}t}+\frac{2\chi}{\lambda_{\operatorname{min}}t}\limsup_{k\to\infty}\frac{\sum_{f(n)\leq k+1}s_{\varepsilon_{n}}}{k}.$
But since the sequence $\{f(n)\}$ can be fixed after the choice of the values $s_{\varepsilon_{n}}$ is made, it simply remains to choose $f$ sparse enough so that $\sum_{f(n)\leq k+1}s_{\varepsilon_{n}}$ grows sublinearly in $k$. Hence, as in the proof for $B=\varnothing$, we get
$\dim{E}_{\infty}\geq\dim H-\frac{4\log\frac{1}{m_{X}(K)}}{\lambda_{\operatorname{min}}t},$
and letting $t\to\infty$ we conclude that the set (5.10) is thick. ∎